OpenStudy (anonymous):

A paper cup is the shape of a cone with a height of 12 cm and radius 4 cm at the top. Water is poured into he cup at a rate of 2cm^3/s. How fast is the water level rising when the water level is 8 cm deep?

OpenStudy (anonymous):

V = pi r^2 h / 3 dV = (pi/3)*(2rdr + r^2 * dh) since h/12 = r/4 r =h/3 and dr =dh/3 dV/dt =(pi/3)*(2(h/3)((dh/dt)/3) + (h/3)^2 * dh/dt) dV/dt = 2 and h = 8 2 =(pi/3)*(2(8/3)((dh/dt)/3) + (8/3)^2 * dh/dt) solve for dh/dt

OpenStudy (anonymous):

maybe I'm working it out wrong but the sheet says that I should get 9/32pi for the answer and instead I'm getting 27/56pi